# Query-Efficient Zeroth-Order Algorithms for Nonconvex Optimization

### v1

@@ D.1  Problem Formulation
Figure 3: 141-bus distribution network.
Mathematically, denote $x\in\mathbb{R}^{d_{x}}$ as the power load of multiple users. Let $D$ denote the reference load received from the grid operator, which imposes a constraint on the distribution network’s net power exchange with the main grid. The power injection of the distribution network is not simply the sum of users’ loads but is determined by nonlinear power flow dynamics. Denote the dynamics as a function of the load levels of multiple users $p_{c}(x):\mathbb{R}^{d_{x}}\to\mathbb{R}$. In our formulation, $p_{c}(x)$ is viewed as a black box, provided that the topology and parameters of the distribution network are unattainable. That means we can only observe the total power consumption $p_{c}(x)$ of a distribution network given the power load $x$ of users.
We apply the 141-bus distribution network model as the nonconvex black box (shown in Figure 3) (Khodr et al.,, 2008). We consider the following problem with $d_{x}=168$:
$\min_{x\in\mathbb{R}^{168}}h(x)=\sum_{i\in[168]}c_{i}(x(i))+\rho(x),$ (32)
$s.t.\;p_{c}(x)\leq D,\;x\in\mathcal{X}=[\underline{x},\overline{x}],$
@@ D.2  Parameters Setting
In our numerical simulation, $\underline{x}$ is also set to 0, and $\overline{x}$ is the nominated load level from the original 141-bus system. The coefficients of cost functions, $a_{i}$ and $b_{i}$, are randomly sampled from the intervals $(0.5,1.5)$ and $(0,5)$, respectively. For the voltage penalty, we set $\underline{v}=0.96\;p.u.$ and $\overline{v}=1.04\;p.u.$. The total load to be curtailed is set to 0.15p.u.=1500kW0.15\;p.u.=1500\;kW. In addition to testing our proposed algorithms, we also compare them with three other algorithms, ZO-MinMax (Liu et al., 2020b, ), ZO-AGP (Xu et al.,, 2024), and stochastic zeroth-order constraint extrapolation (SZO-ConEX) (Nguyen and Balasubramanian,, 2023). For all tests, we compute $\|\mathfrak{g}(x_{k},y_{k})\|$ as the stationarity measures. The constraint violation is measured by $p_{c}(x)-D$. In Table 2, the relative error is computed by $(h(x_{k})-h^{*})/h^{*}$ based on the optimal objective function value $h^{*}=0.3841$.
@@
For ZOB-GDA, we consider four scenarios with batch sizes $b=1,10,50,168$, where the step sizes are set as $\alpha=0.03,0.025,0.005,0.00035$ and $\beta=0.01\alpha$, respectively. For ZOB-SGDA, we set $p=10$ and $\gamma=0.3$, and adopt the same batch-size scenarios and corresponding step sizes as ZOB-GDA. For the benchmark algorithms, ZO-MinMax is implemented with $\alpha=\beta=5\times 10^{-6}$; since it involves constraint-handling techniques, we adopt a decaying penalty parameter $\delta_{k}=\min(50/k,0.1)$. For SZO-ConEx, we set $\alpha=\beta=5\times 10^{-6}$. For ZOAGP, we choose $\alpha=\beta=0.00035$. We set the maximum iteration steps as $K=20000$. For the smoothing radius, ZOB-GDA, ZOB-SGDA, and ZOAGP use $r_{k}=\min(10^{-1}/(k^{1.2}),2\times 10^{-4})$, while SZO-ConEx and ZO-MinMax use $r_{k}=\min(10^{-2}/(k+4000)^{1.1},1\times 10^{-5})$. All parameters are selected for the best performance among multiple options. All experiments are conducted on a MacBook Pro laptop equipped with an Apple M1 Pro SoC (10-core CPU: 8 performance cores and 2 efficiency cores) and 16 GB of unified memory.
D.3  Numerical Results for Noisy Cases
In this subsection, we further test our proposed algorithms, ZOB-GDA and ZOB-SGDA, under noisy observations to validate their robustness. The observed values of $p_{c}(x)$ are perturbed by additive Gaussian noise with zero mean and a standard deviation of $5$ kW. In noisy cases, the smoothing radius is set as $r_{k}=\min(400/(k^{1.2}),4\times 10^{-3})$, while all the other parameter settings remain unchanged as in the noise-free cases. Each algorithm is tested with 20 runs, and the average performance is presented in Figure 4. The average number of queries required to generate solutions with different qualities is summarized in Table 3. The results show that our algorithms also exhibit satisfactory performance in noisy cases without degrading significantly, which demonstrates the robustness of our methods.
Figure 4: Performance of ZOB-GDA and ZOB-SGDA under noisy observations. (a), (b), and (c) present the objective function value, constraint violation, and stationarity measure of ZOB-GDA. (d), (e), and (f) present the corresponding results for ZOB-SGDA.
Table 3: Average numbers of iterations and queries required to generate solutions with certain levels of relative errors and zero constraint violation under noisy observation.
Relative error 10% 1% 0.1%
Iteration Complexity Iteration Complexity Iteration Complexity
ZOBGDA b=1 680.8 1361.6 1263.3 2526.6 1816.75 3633.5
b=10 75.05 825.55 139.25 1531.75 194.55 2140.05
b=50 21.65 1104.15 60.65 3093.15 97.05 4949.55
b=168 51.9 8771.1 194.8 32921.2 255.25 43137.25
ZOB-SGDA b=1 687.4 1374.8 1282.15 2564.3 1937.35 3874.7
b=10 76.95 846.45 141 1551 191.05 2101.55
b=50 21.8 1111.8 61.6 3141.6 101.2 5161.2
b=168 52.25 8830.25 196.65 33233.85 257.25 43475.25

### v2

@@ D.1  Energy Management
Mathematically, denote $x\in\mathbb{R}^{d_{x}}$ as the power load of multiple users. Let $D$ denote the reference load received from the grid operator, which imposes a constraint on the distribution network’s net power exchange with the main grid. The power injection of the distribution network is not simply the sum of users’ loads but is determined by nonlinear power flow dynamics. Denote the dynamics as a function of the load levels of multiple users $p_{c}(x):\mathbb{R}^{d_{x}}\to\mathbb{R}$. In our formulation, $p_{c}(x)$ is viewed as a black box, provided that the topology and parameters of the distribution network are unknown. That means we can only observe the total power consumption $p_{c}(x)$ of a distribution network given the power load $x$ of users.
Figure 4: 141-bus distribution network
We apply the 141-bus distribution network model as the nonconvex black box (shown in Figure 4) (Khodr et al., 2008). We consider the following problem with $d_{x}=168$:
$\min_{x\in\mathbb{R}^{168}}h(x)=\sum_{i\in[168]}c_{i}(x(i))+\rho(x),\quad\text{s.t.}\;p_{c}(x)\leq D,\;x\in\mathcal{X}=[\underline{x},\overline{x}],$ (31)
@@
Parameters Setting. In our numerical simulation, $\underline{x}$ is also set to 0, and $\overline{x}$ is the nominal load level from the original 141-bus system. The coefficients of cost functions, $a_{i}$ and $b_{i}$, are randomly sampled from the intervals $(0.5,1.5)$ and $(0,5)$, respectively. For the voltage penalty, we set $\underline{v}=0.96\;p.u.$ and $\overline{v}=1.04\;p.u.$. The total load to be curtailed is set to 0.15p.u.=1500kW0.15\;p.u.=1500\;kW. In addition to testing our proposed algorithms, we also compare them with three other algorithms, ZO-MinMax, ZOAGP, and stochastic zeroth-order constraint extrapolation (SZO-ConEx). For all tests, we compute $\|\mathfrak{g}(x_{k},y_{k})\|$ as the stationarity measure. The constraint violation is measured by $p_{c}(x)-D$. In Table 2, the relative error is computed by $(h(x_{k})-h^{*})/h^{*}$ based on the optimal objective function value $h^{*}=0.3841$.
@@
For ZOB-GDA, we consider four scenarios with batch sizes $b=1,10,50,168$, where the step sizes are set as $\alpha=0.03,0.015,0.005,0.00035$ and $\beta=0.01\alpha$, respectively. For ZOB-SGDA, we set $p=10$ and $\gamma=0.3$, and adopt the same batch-size scenarios and corresponding step sizes are set as $\alpha=0.03,0.025,0.005,0.00035$ and $\beta=0.01\alpha$. For the benchmark algorithms, ZO-MinMax is implemented with $\alpha=\beta=5\times 10^{-6}$; since it involves constraint-handling techniques, we adopt a decaying penalty parameter $\delta_{k}=\min(50/k,0.1)$. For SZO-ConEx, we set $\alpha=\beta=5\times 10^{-6}$. For ZOAGP, we choose $\alpha=\beta=0.00035$. We set the maximum iteration steps as $K=20000$. For the smoothing radius, ZOB-GDA, ZOB-SGDA, and ZOAGP use $r_{k}=\min(10^{-1}/k^{1.2},2\times 10^{-4})$, while SZO-ConEx and ZO-MinMax use $r_{k}=\min(10^{-2}/(k+4000)^{1.1},1\times 10^{-5})$. All parameters are selected for the best performance among multiple options. All experiments are conducted on a MacBook Pro laptop equipped with an Apple M1 Pro SoC (10-core CPU: 8 performance cores and 2 efficiency cores) and 16 GB of unified memory.
D.2  High-Dimensional Parameter Optimization
In the 1000-dimensional constrained nonconvex optimization problem, each algorithm is tested 20 times.
Parameter Settings The matrix $B\in\mathbb{R}^{1000\times 1000}$ and the constraint vector $q\in\mathbb{R}^{1000}$ are generated randomly with i.i.d. entries following $\mathcal{N}(0,\sigma^{2})$ with $\sigma=1/10\sqrt{10}$, and remain fixed throughout each run. We set $\bar{c}=0.5$ for all tests. For ZOB-GDA, we apply a tuned block size $b=30$, and the step sizes are set as $\alpha=0.5$ and $\beta=0.5$. For ZOB-SGDA, we set $p=0.1$, $\gamma=0.9$, $\alpha=0.5$ and $\beta=5$. For the comparison algorithms, ZO-MinMax is implemented with $\alpha=\beta=0.2$. For SZO-ConEx, we set $\alpha=0.3$ and $\beta=0.003$. For ZOAGP, we set $\alpha=0.1$ and $\beta=5$. The maximum budget of function evaluations is set as $1\times 10^{6}$. For the smoothing radius, ZOB-GDA and ZOB-SGDA use $r_{k}=\min(10^{-2}/k,1\times 10^{-3})$; ZO-MinMax and SZO-ConEx use $r_{k}=\min(10^{-2}/k,2\times 10^{-4})$; ZOAGP uses $r_{k}=\min(10^{-3}/k,1\times 10^{-4})$. All parameters are selected for the best performance among multiple options.